hi, im working on complements and suplements of angles and i had a question from my worksheet.
"The measure of the complement of an angle is 45+x. The measure of the supplement of an angle is 15x-47. What is the measure of the angle?"
My work: 45 + x = 90 15x - 47 = 180 15x 227
-45 = -45 +47 = +47 ---- = --- =(x = 15.13333333)
------------- --------------- 15x 15x
(x = 45) 15x = 227
what i dont get is what to do next. the question is asking for only one of the angles but i have two x's. If i plugg them in then i have two different answers. What should i do next to finish this problem?

Answer by stanbon(75887)

The measure of the complement of an angle is 45+x. The measure of the supplement of an angle is 15x-47. What is the measure of the angle?
--------------
Let the measure of the angle be "y".
EQUATIONS:
y + 45+x = 90
y + 15x-47 = 180
----------------
Rearrange to get:
y + x = 45
y + 15x = 227
-------------
Subtract 1st from 2nd to solve for "x":
14x = 182
x = 13
--------------------
Substitute to solve for "y":
y + 13 = 45
y = 32 (final answer)
=======================
Cheers,
Stan H.